Total 4 marks estion 6 )Explain the difference between class limit and class bounda ) Use the frequency table, Table 1, overleaf to answer the following questions [2 marks) Table 1 0-9 10-19 20-29 30-39 (1) What is the class width? (ii) What is the sample size? (ii) What is the lower class limit of the first class? 14 ivhat is the lowerlass boundary of the first dis? 4 mark continued ov Page 20 of Total
Answer #1

a) Class limit: It is the smallest and the largest value in a class , often called as Lower limit and upper limit in a single class.

Whereas,

Class boundary: It is the midpoint value between upper class limit of one class and lower class limit of next class.

b) i) Class width= It is the difference between the upper class limit or lower class limit of consecutive classes, hence

10-0=10 or 19-9=10 is the class width.

ii) Sample size = It is sum of the all the class frequency, ie 3+14+11+5=33

iii) Class limit as explained in part a) here lower class limit of class 1 is 0.

iv) The lower class boundry for first class would be -0.95